A real estate developer is considering investing in a shopping mall on the outskirts of Dallas, Texas. Various parcels of land are being evaluated. Of particular importance is the income in the area surrounding the proposed mall. A random sample of families is selected near each proposed mall. Following are the sample results. At the 0.05 significance level, can the developer conclude there is a difference in the mean income? Use the usual 5 step hypothesis testing procedure.
McKinney Richardson Allen Plano
4 2 1 2
8 3 2 3
4 4 3 1
3
a) State the null and alternate hypothesis
b) What is alpha?
c) What is n?
d) What is K?
e) Formulas:
a) SS total=
b) SSE=
c) SST=
d) dfn=
e) dfd=
f) MST=
g) MSE=
h) F=

Thanks in Advance! :-) Also, the numbers in the cities column are listed individualy and vertically. For some reason they get bunched up on the system.

Answer by stanbon(75887)

A random sample of families is selected near each proposed mall. Following are the sample results. At the 0.05 significance level, can the developer conclude there is a difference in the mean income? Use the usual 5 step hypothesis testing procedure.
McKinney Richardson Allen Plano
...4........ 2...... 1..... 2
...8........ 3...... 2..... 3
...4........ 4...... 3..... 1
...3
a) State the null and alternate hypothesis
Ho: the means are the same
Ha: at least one of the means is different
-----------------------
b) What is alpha? 5%
------------------------
c) What is n? n is number of samples within a category = 3
-------------------
d) What is K? k is number of categories = 4
-------------------
e) Formulas:
a) SS total= summation of (x-mean of sample means)^2 =18.1731...
b) SSE= summation of (ni -1)*xi^2
c) SST=
d) dfn= numerator degrees of freedom = k-1 = 3
e) dfd= denominator degrees of freedom = N-k = 1
f) MST=[SS(total)]/[N-1]
g) MSE=[SS(error)]/(N-k)
h) F= variance between samples/variance within sample = 2.62743...
-------------------
p-value = 0.11427...
Conclusion:
Fail to reject Ho.
======================
Cheers,
Stan H.